# Breaking the Script Barrier in Multilingual Pre-Trained Language Models with Transliteration-Based Post-Training Alignment

### v1

@@ 6 Conclusion
In this work, we propose a transliteration-based post-training method that contains both sequence and token-level objectives to improve the cross-lingual/script alignment of mPLMs and thus boost their zero-shot cross-lingual transfer performance. We apply our post-training method to fine-tune Glo500 on two language groups that share areal features and have extensive lexical overlap. Our extensive experiments using different source languages show that our aligned models consistently outperform the original Glot500 model. In particular, our method enhances the alignment between related languages and, therefore, improves cross-lingual transfer between these languages. We also analyze the impact of different training objectives and show that the sequence and token-level alignment objectives are both critical for achieving the best performance across different tasks.
@@
Even though the mPLM is fine-tuned with our method, where the transliterated text is used as an auxiliary input, the mPLM has only seen the Latin transliterations during its pre-training phase. This can limit the performance of the post-trained aligned models, especially for languages with complex scripts. An extension of this work could expand the vocabulary to include the subwords from Latin transliterations as done by Liu et al. (2024b) before fine-tuning or continued-pre-train the model on the transliterated text so that the models can be more effective in modeling the transliterated data.
@@ Acknowledgments
This research was supported by DFG (grant SCHU 2246/14-1) and The European Research Council (NonSequeToR, grant #740516). We would like to thank Leibniz Supercomputing Center (LRZ) and Munich Center for Machine Learning (MCML) for providing the computational resources.
@@ Appendix B Downstream Tasks Fine-Tuning
For the downstream tasks that require fine-tuning, we further fine-tune the post-trained aligned models on the training set of a given source language, select the best checkpoint with early stopping based on the f1 score on the source language’s validation set, and evaluate the macro F1 score on the test sets of the remaining languages in each group. The fine-tuning results are averaged over five different seeds. All models are fine-tuned on a single GPU. Unless otherwise stated, the same optimizer and scheduler settings are used as in the post-training alignment phase.
@@
We evaluate the models on the Universal Dependencies (UD) v2.11 part-of-speech tagging dataset (de Marneffe et al., 2021), a multilingual dataset consisting of sentences annotated with 17 universal POS tags. We use the same fine-tuning setup as the NER task, except for number of epochs, which is set to 10.

### v2

@@ 6 Conclusion
In this work, we propose a transliteration-based post-training method that includes both sequence and token-level objectives to improve the cross-lingual alignment of mPLMs and thus boost their cross-lingual transfer performance. We apply our post-training method to fine-tune Glot500 on two language groups that share areal features and have extensive lexical overlap. Our extensive experiments using different source languages show that our aligned models consistently outperform the original Glot500 model. In particular, our method enhances the alignment and cross-lingual transfer between related languages. We also analyze the impact of different training objectives and show that the sequence and token-level alignment objectives are both critical for achieving the best performance across different tasks.
@@
Even though the mPLM is fine-tuned with our method, where the transliterated text is used as an auxiliary input, the mPLM has only seen the Latin transliterations during its pre-training phase. This can limit the performance of the post-trained aligned models, especially for languages with complex scripts. An extension of this work could expand the vocabulary to include the subwords from Latin transliterations as done by Liu et al. (2024b) before fine-tuning or continually pre-training the model on the transliterated text so that the models can be more effective in modeling the transliterated data.
@@
Future work could also explore explicit token-level alignment objectives, such as minimizing the L2 distance between word/subword pair representations as done by Cao et al. (2020). However, this kind of alignment cannot be directly applied to many languages in our study. For the South+East Asian Languages group, most language scripts are logographic or do not use word boundaries, making it challenging to provide an explicit supervision signal for token-level alignment, as much of this information is lost during the transliteration process.
@@ Acknowledgments
This work was funded by Deutsche Forschungsgemeinschaft (project SCHU 2246/14-1) and The European Research Council (NonSequeToR, grant #740516). We would like to thank Leibniz Supercomputing Center (LRZ) and Munich Center for Machine Learning (MCML) for providing the computational resources.
@@ Appendix B Downstream Tasks Fine-Tuning
For the downstream tasks that require fine-tuning, we further fine-tune the aligned models on the training set of a given source language, select the best checkpoint with early stopping based on the f1 score on the source language’s validation set, and evaluate the macro F1 score on the test sets of the remaining languages in each group. The results are averaged over five different seeds. All models are fine-tuned on a single GPU. Unless otherwise stated, the same optimizer and scheduler settings are used as in the post-training alignment phase.
@@
We evaluate the models on the Universal Dependencies (UD) v2.11 part-of-speech tagging dataset (de Marneffe et al., 2021), a multilingual dataset consisting of sentences annotated with 17 universal POS tags. We use the same fine-tuning setup as the NER task, except for the number of epochs, which is set to 10.